OpenStudy (anonymous):

A right triangle has one side of length 7 and a hypotenuse of length 10. How long is the third side of the triangle?

OpenStudy (anonymous):

use Pythagorus theorem. Do you know its formula?

OpenStudy (anonymous):

nope can you give it to me please

OpenStudy (anonymous):

Let c be the hypotenuse. \[c^2=a^2+b^2\] now, substitute c=10 and a=7 \[10^2=7^2+b^2\] \[100-49=b^2\] \[\sqrt51=b\]

OpenStudy (anonymous):

your length would be \[\sqrt{51}\] . Depending on your answer, you might want to simplify it further.

OpenStudy (anonymous):

http://www.mathsisfun.com/pythagoras.html here's a link to check out.

OpenStudy (anonymous):

if you didn't quite get my explanation.

OpenStudy (anonymous):

Yep that is what i got too

OpenStudy (anonymous):

thank you very much